# Continuous-time Data-driven Barrier Certificate Synthesis

### v1

@@ Abstract
We consider the problem of verifying safety for continuous-time dynamical systems. Developing upon recent advancements in data-driven verification, we use only a finite number of sampled trajectories to learn a barrier certificate, namely a function which verifies safety. We train a safety-informed neural network to act as this certificate, with an appropriately designed loss function to encompass the safety conditions. In addition, we provide probabilistic generalisation guarantees from discrete samples of continuous trajectories, to unseen continuous ones. Numerical investigations demonstrate the efficacy of our approach and contrast it with related results in the literature. For the purpose of Open Access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.
1 Introduction
Ensuring the safety of Continuous-time dynamical systems is of critical importance in an increasingly autonomous world [8, 15, 19]. As it is often infeasible to model system behaviour precisely, and making direct use of system data to verify behaviour is of interest [1, 12].
@@ 1 Introduction
A technique to verify properties of dynamical systems involves discretising the state space [3], under approximation guarantees, and verifying the resulting model. Alternatively, the use of certificates [2, 18] allows one to analyse directly the continuous-state system. These certificates map system’s states to real values, and exhibit certain properties that are relevant for analysis: here in particular we construct safety certificates for continuous-time systems, but extensions to more complex certificates are possible  [19].
@@
There are a number of techniques for synthesising such certificates. In the case that an exact model is known, one can use a polynomial function as a certificate to formulate a convex sum-of-squares problem [17]. Recent work in this area investigated the use of neural networks as certificates [8], which represent a class of general function approximators.
@@
When obtaining an exact model is infeasible, we turn to data-driven techniques. One method for employing data in certificate synthesis is through the use of state pairs (i.e. states, and next-states), sampled from across the domain of interest. Such techniques are investigated in [15] for deterministic systems, and in [21] for stochastic systems. Both these works make use of the techniques in [11, 14] to bound the distance between what is referred to as a robust program, and its sample based counterpart. As discussed in [14, Remark 3.9], such techniques exhibit an exponential growth in the dimension of the sampling space (here the state space), and also require access to the underlying probability distribution.
@@
Alternatively, one can consider using entire trajectories as samples, hence only requiring access to realistic runs of the system. This technique is explored in [19] for discrete-time systems. This work leverages the so-called scenario approach [5, 10] and the notion of compression [13, 7], and provides a constructive algorithm for the general pick-to-learn framework [16], to provide a probably approximately correct (PAC) bound on the correctness of certificates for newly sampled trajectories. These techniques are restricted to discrete-time systems, since they perform calculations on entire trajectories, which is infeasible in continuous-time. Hence, here we develop techniques for continuous-time systems by analysing samples from complete trajectories and bounding the difference between the safety of these approximations and the complete trajectories.
Our main contribution is the synthesis of neural barrier certificates for continuous-time dynamical systems, accompanied by PAC generalisation guarantees. These certificates are useful per se as a proxy for safety, and open the road for their use in control synthesis, an area of ongoing research.
@@
Notation. We use $\{\xi_{k}\}_{k=0}^{K}$ to denote a sequence indexed by $k\in\{0,1,\dots,K\}$. $B\models\psi$ defines condition satisfaction i.e., it evaluates to true if the quantity $B$ on the left satisfies the condition $\psi$ on the right, e.g., $x=1\models x>0$ evaluates to true and $x=-1\models x>0$ evaluates to false. Using ⊧$\not\models$̸ represents the logical inverse of this (i.e., condition dissatisfaction). By $(\forall\xi\in\Xi)B\models\psi(\xi)$ we mean that some quantity $B$ satisfies a condition $\psi$ which, in turn, depends on some parameter $\xi$, for all $\xi\in\Xi$. We use $\xi_{[0,k]}$ to refer to a (possibly infinite) subsequence $\{\xi_{0},\dots,\xi_{k}\}$ of a sequence.
2 Certificates
In this work, we focus on barrier certificates, but emphasise that our techniques naturally extend to the more complex certificates in [19]. To this end, we begin by defining a dynamical system, before considering the barrier certificate and associated safety property it verifies.
2.1 Continuous-Time Dynamical Systems
@@
We consider a bounded state space $X\subseteq\mathbb{R}^{n}$, and a dynamical system whose evolution starts at an initial state $x(0)\in X_{I}$, where $X_{I}\subseteq X$ denotes the set of possible initial conditions. From an initial state, we uncover a finite trajectory, i.e., a continuous sequence of states $\xi=\{x(t)\}_{t\in[0,T]}$, where $T\in\mathbb{R}$ and $x(t)\colon t\rightarrow\mathbb{R}^{n}$, by following the dynamics
@@
We only require $f\colon X\rightarrow\mathbb{R}^{n}$ to be Lipschitz continuous, thus enforcing unique solutions. The set of all possible trajectories $\Xi\subseteq X_{I}\times X^{(0,T]}$ is then the set of all trajectories starting from the initial set $X_{I}$.
In Section 3, we discuss how to use a finite set of trajectories to build safety certificates, and accompany them with generalization guarantees with respect to their validity for any trajectory. To allow for the synthesis of such certificate, as discussed in Section 4, we must store and perform calculations on such trajectories that play the role of samples. For this reason, we consider a finite $T$, and take samples along the continuous trajectory. Hence, we define the time-discretised trajectory
@@
2.2 Safety Property
We use $\phi(\xi)$ to refer to the safety property under study, evaluated on a trajectory $\xi\in\Xi$, defined as follows.
@@ 2.2 Safety Property
Consider (1), and let $X_{I},X_{U}\subset X$ with $X_{I}\cap X_{U}=\varnothing$ denote an initial and an unsafe set, respectively. If, for all $\xi\in\Xi$,
$\phi(\xi)\vcentcolon=\forall t\in[0,T],x(t)\notin X_{U},$
holds, then we say that $\phi$ encodes a safety property. $\Xi$ denotes the set of trajectories consistent with (1) and with initial state contained within $X_{I}$.
By the definition of $\phi$, it follows that verifying a system exhibits the safety property is equivalent to verifying all trajectories emanating from the initial set avoid the unsafe set for all time instances, until horizon $T$.
2.3 Barrier Certificate
We now define the relevant criteria necessary for a certificate $B$ to verify a safety property. We fix a time horizon $T<\infty$ and assume that $B$ is continuous, so that when considering the supremum/infimum of $B$ over $X$ (already assumed to be bounded) or over some of its subsets, this is well-defined. Consider:
@@ 2.3 Barrier Certificate
$\evaluated{\frac{dB}{dt}}_{x\in\xi}<\frac{1}{T}\Big{(}\inf_{x\in X_{U}}B(x)-\sup_{x\in X_{I}}B(x)\Big{)},$ (5)
where we may expand the Lie derivative as
$\frac{dB}{dt}=\frac{\partial B}{\partial x}\frac{dx}{dt}=\frac{\partial B}{\partial x}f(x),$ (6)
and hence recognise that this depends on the system dynamics $f(x)$. We place the following assumption on this Lie derivative, necessary for our analysis later.
Assumption 1 (Lie Derivative is Lipschitz)
@@
Notice that even if $\inf_{x\in X_{U}}B(x)-\sup_{x\in X_{I}}B(x)>0$, i.e., in the case where the last condition encodes an increase of $B$ along the system trajectories, the system still avoids entering the unsafe set. This is established in the proof of Proposition 8 below.
A graphical representation of these conditions is found in Figure 3. The level set (with dashed line) is the set obtained when the certificate value is 00. The decrease condition then means that we never leave the sublevel set.
@@
i.e. the maximum value along a trajectory is less than the infimum over the unsafe region and hence $x(t)\notin X_{U},t=[0,T]$ (notice that $x(0)\notin X_{U}$ holds since $X_{I}\cap X_{U}=\varnothing$). The latter implies that all trajectories that start in $X_{I}$ avoid entering the unsafe set $X_{U}$, thus concluding the proof.
To synthesise one of these deterministic certificates, we require complete knowledge of the behaviour $f$ of the dynamical system, to allow us to evaluate the Lie derivative $\frac{dB}{dt}$. This may be impractical, and we therefore use data-driven techniques to learn a certificate.
3 Data-Driven Certificates
@@
To obtain our sample set, we consider $N$ independent and identically distributed (i.i.d.) initial conditions, sampled according to probability distribution $\mathbb{P}$, namely $\{x^{i}(0)\}_{i=1}^{N}\sim\mathbb{P}^{N}$. Initializing the dynamics from each of these initial states, we unravel a set of continuous-time trajectories $\{\xi^{i}\}_{i=1}^{N}$. Since there is no stochasticity in the dynamics, we can equivalently say that trajectories (generated from the random initial conditions) are distributed according to the same probabilistic law; hence, with a slight abuse of notation, we write $\xi\sim\mathbb{P}$. We impose the following assumption.
Assumption 1 (Non-concentrated Mass).
@@
3.1 Problem Statement
Since we are now dealing with a sample-based problem, we will construct probabilistic certificates and hence probabilistic guarantees on the satisfaction of a given property.
Denote by $B_{N}$ a barrier certificate, we introduce the subscript $N$ to emphasize that this certificate is constructed on the basis of time-discretised sampled trajectories $\{\tilde{\xi}^{i}\}_{i=1}^{N}$.
Problem 2 (Probabilistic Property Guarantee).
Consider $N$ sampled trajectories, and fix a confidence level $\beta\in(0,1)$. We seek a property violation level $\epsilon\in(0,1)$ such that
$\mathbb{P}^{N}$$\big{\{}\{\tilde{\xi}^{i}\}_{i=1}^{N}\in\tilde{\Xi}^{N}:~{}$
$\mathbb{P}\{\xi\in\Xi\colon B_{N}\not\models\psi^{s}\wedge\psi^{\Delta}(\xi)\}\leq\varepsilon\big{\}}\geq 1-\beta.$ (10)
In words, finding a solution to Problem 2 requires determining an $\epsilon\in(0,1)$, such that with confidence at least $1-\beta$, the probability that $B_{N}$ does not satisfy the condition $\psi^{s}\wedge\psi^{\Delta}(\xi)$ for another sampled continuous-time trajectory $\xi\in\Xi$ is at most equal to that $\epsilon$. As such, with a certain confidence, a certificate $B_{N}$ trained on the basis of $N$ sampled trajectories, will remain a valid certificate with probability at least $1-\epsilon$. Note that the outer probability in (10) (as well as in similar statements below) refers to the selection of discretized trajectories, $\{\tilde{\xi}^{i}\}_{i=1}^{N}$ (as these are used for training), while the inner probability refers to the selection of a continuous time one $\xi$, as this captures the desired generalization properties.
3.2 Probabilistic Guarantees
Consider a mapping $\mathcal{A}$ such that $B_{N}=\mathcal{A}(\{\xi^{i}\}_{i=1}^{N})$ as an algorithm that, based on $N$ samples, returns a certificate $B_{N}$. We call as compression set of such an algorithm any subset of the input that returns the same certificate. That is, a sample-subset $\mathcal{C}_{N}\subseteq\{\xi^{i}\}_{i=1}^{N}$ is a compression set if $\mathcal{A}(\mathcal{C}_{N})=\mathcal{A}(\{\xi^{i}\}_{i=1}^{N})$. In Algorithm 1, we provide a specific synthesis procedure through which $\mathcal{A}$ (and hence the certificate $B_{N}$) can be constructed. This algorithm is adapted from [19], with appropriate modifications to allow for continuous-time dynamics; we discuss this in the next section. Using the results of [19], and if we have access to samples of the time-derivative $f(x)$, we can provide a guarantee over the time-discretised trajectories, as stated next.
Theorem 3 (Probabilistic Guarantees [19]).
Consider Assumption 1, and let $B_{N}$ and $\mathcal{C}_{N}$ be the certificate and compression set, respectively, returned by Algorithm 1. Fix $\beta\in(0,1)$, and for $k<N$, let let $\varepsilon(k,\beta,N)$ be the (unique) solution to the polynomial equation in the interval $[k/N,1]$
@@ 3.2 Probabilistic Guarantees
Unfortunately, this does not provide us with a guarantee on the property satisfaction for the continuous-time trajectories, i.e. our trajectory may violate the safety property between sampled states $x(t_{i})$ and $x(t_{i+1})$, as the inner probability refers to a choice of $\tilde{\xi}$ as opposed to $\xi$.
If the Lipschitz continuity in Assumption 1 holds, we can additionally offer guarantees on newly sampled continuous-time trajectories, based only on time-discretised trajectories, and approximate derivatives. In this case, we require a tightening of the derivative condition in (5) by some value $d\in\mathbb{R}$, to ensure that the decrease condition is maintained between sample times. Denote this condition (over discretized trajectories) as $\psi^{\Delta}_{d}(\tilde{\xi})$, defined by the inequality
@@
$<\frac{1}{T}\Big{(}\inf_{x\in X_{U}}B(x)-\sup_{x\in X_{I}}B(x)\Big{)}-d.$
Define by $L_{B}$ and $L_{f}$ the Lipschitz constants of the certificate derivative $\frac{\partial B_{\theta^{\star}}}{\partial x}$ and of the dynamics $f(x)$ respectively, and by $\mathcal{M}_{B},\mathcal{M}_{f}$ bounds on their norms, namely $\sup_{x}\|\frac{\partial B_{\theta^{\star}}}{\partial x}\|\leq\mathcal{M}_{B}$ and $\sup_{x}\|f(x)\|\leq\mathcal{M}_{f}$. Then, the value $d$ is defined as
$d=\overline{t}\mathcal{M}_{f}\left(\mathcal{M}_{B}\mathcal{L}_{f}+\mathcal{M}_{f}\mathcal{L}_{B}\right),$ (14)
where $\overline{t}=\max_{k=1,\dots,M}(t_{k}-t_{k-1})$.
Theorem 4 (Continuous-Time Guarantees).
@@
The proof of this is achieved by bounding the difference between the safety of the continuous-time trajectories and their time-discretised approximations, and is in the Appendix.
4 Certificate Synthesis
In order to learn a barrier certificate from samples, we consider a neural network, a well-studied class of function approximators that generalize well to a given task. Denote all tunable neural network parameters by a vector $\theta$. We then have that our certificate $B_{N}$ depends on $\theta$. For the results of this section, we simply write $B_{\theta}$ and drop the dependency on $N$ to ease notation.

### v2

@@ Abstract
We consider the problem of verifying safety for continuous-time dynamical systems. Developing upon recent advancements in data-driven verification, we use only a finite number of sampled trajectories to learn a barrier certificate, namely a function which verifies safety. We train a safety-informed neural network to act as this certificate, with an appropriately designed loss function to encompass the safety conditions. In addition, we provide probabilistic generalisation guarantees from discrete samples of continuous trajectories, to unseen continuous ones. Numerical results demonstrate the efficacy of our approach and contrast it with related results in the literature. For the purpose of Open Access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.
I Introduction
@@
Ensuring the safety of continuous-time dynamical systems is of critical importance in an increasingly autonomous world [1, 2, 3]. It is often infeasible to model system behaviour precisely, thus making direct use of system data to verify behaviour is of interest [4, 5]. A technique for verifying properties of dynamical systems involves discretising the state space [6], under approximation guarantees, and verifying the resulting model. Alternatively, the use of certificates [7, 8] allows one to directly analyse the continuous-state system. These certificates map the system’s states to real values, and exhibit certain properties that are relevant for analysis: here, in particular, we construct safety certificates for continuous-time systems, but extensions to the more complex certificates in [3] are possible.
@@
There are a number of techniques for synthesising such certificates. In the case that an exact model is known, one can use a polynomial function as a certificate to formulate a convex sum-of-squares problem [9]. Recent work in this area investigated the use of neural networks (which represent a class of general function approximators) as certificates [1].
@@
Obtaining a model of the system, however, is generally a difficult task. It requires domain-specific knowledge and, since parts of the system may often not be known exactly, such a model will only constitute an approximation. To alleviate these issues, we turn to model-free data-driven techniques. One method for employing data in certificate synthesis is through the use of state pairs (i.e. states, and next-states), sampled from across the domain of interest. Such techniques are investigated in [2] for deterministic systems, and in [10] for stochastic systems. Both these works make use of the techniques in [11, 12] to bound the distance between what is referred to as a robust program, and its sample-based counterpart. As discussed in [12, Remark 3.9], such techniques exhibit an exponential growth in the dimension of the sampling space (here the state space), and also require access to the underlying probability distribution.
@@
Alternatively, one can consider using entire trajectories as samples, hence only requiring access to runs compatible with the dynamics of the system. This is explored in [3] for discrete-time systems. This work leverages the so-called scenario approach [13, 14] and the notion of compression [15, 16], and provides a constructive algorithm for the general pick-to-learn framework [17], to provide a probably approximately correct (PAC) bound on the correctness of certificates for newly sampled discrete trajectories. However, these guarantees are no longer valid when it comes to continuous-time systems. Here, we extend these developments to allow constructing a safety/barrier certificate based on a finite number of discretised trajectories of the continuous-time system; however, we provide guarantees on the safety of a new unseen continuous-time trajectory. The latter does not follow from the results in [3].
Our contributions can be summarised as follows:
  • We develop novel theoretical results that extend the results of [3] (which was limited in scope to discrete-time systems) to verify entire continuous trajectories, using discretised approximations for computations;
  • We complement the results in [2], following an alternative approach for data-driven certificate synthesis which, however, does not scale exponentially with respect to the dimension of the underlying state space;
  • Our approach is entirely data-driven, avoiding the use of satisfiability modulo theories (SMT) solvers, which are computationally expensive and require a system model.
@@
Notation. We use $\{\xi_{k}\}_{k=0}^{K}$ to denote a sequence indexed by $k\in\{0,1,\dots,K\}$. $B\models\psi$ defines condition satisfaction, i.e., it evaluates to true if the quantity $B$ on the left satisfies the condition $\psi$ on the right. Using $\not\models$ represents the logical inverse of this (i.e., condition dissatisfaction). By $(\forall\xi\in\Xi)B\models\psi(\xi)$ we mean that some quantity $B$ satisfies a condition $\psi$ which, in turn, depends on some parameter $\xi$, for all $\xi\in\Xi$. We use $\xi_{[0,k]}$ to refer to a subsequence $\{\xi_{0},\dots,\xi_{k}\}$ of a sequence.
II Safety and Barrier Certificates
@@
We focus on barrier certificates as the tool to verify safety; however, our techniques naturally extend to more complex certificates, as in [3]. Consider a bounded state space $X\subseteq\mathbb{R}^{n}$, and a continuous-time dynamical system whose evolution starts at an initial state $x(0)\in X_{I}$, where $X_{I}\subseteq X$ denotes the set of possible initial conditions. From an initial state, we unravel a finite trajectory, i.e., a continuous sequence of states $\xi=\{x(t)\}_{t\in[0,T]}$, where $T\in\mathbb{R}$ and $x(t)\colon[0,T]\rightarrow\mathbb{R}^{n}$, by following the dynamics
@@
We only require $f\colon X\rightarrow\mathbb{R}^{n}$ to be Lipschitz continuous, thus allowing for existence and uniqueness of solutions. The set of all possible trajectories $\Xi\subseteq X_{I}\times X^{(0,T]}$ is then the set of all trajectories starting from the initial set $X_{I}$.
In Section III, we discuss how to use a finite set of trajectories to build safety certificates, and accompany them with generalisation guarantees with respect to their validity for a new, unseen trajectory. We will also consider time-discretised versions of continuous-time trajectories. To this end, define the time-discretised trajectory
@@
Consider (1), and let $X_{I},X_{U}\subset X$ with $X_{I}\cap X_{U}=\varnothing$ denote an initial and an unsafe set, respectively. We say that $\phi$ encodes a safety property for a trajectory $\xi\in\Xi$ if,
$\phi(\xi)\vcentcolon=\forall t\in[0,T],x(t)\notin X_{U}.$
By the definition of $\phi$, it follows that verifying that a trajectory exhibits the safety property is equivalent to verifying that a trajectory emanating from the initial set avoids the unsafe set for all time instances, until the horizon $T<\infty$.
We now define the relevant criteria for a certificate $B$ to verify a safety property. Assume that $B$ is continuous, so that when considering the supremum/infimum of $B$ over $X$ (already assumed to be bounded) or over some of its subsets, this is well-defined. Consider:
@@
[…]
where $\frac{dB}{dt}=\frac{\partial B}{\partial x}f(x)$, and hence recognise that this depends on the system dynamics $f(x)$.
Assumption 1 (Lipschitz Derivative)
@@
Note that even if $\inf_{x\in X_{U}}B(x)-\sup_{x\in X_{I}}B(x)>0$, i.e., if the last condition encodes an increase of $B$ along the system trajectories, the system still avoids entering the unsafe set. This is established in the proof of Proposition 1 below.
@@
i.e. the maximum value along a trajectory is less than the infimum over the unsafe region and hence $x(t)\notin X_{U},t=[0,T]$ (notice that $x(0)\notin X_{U}$ holds since $X_{I}\cap X_{U}=\varnothing$). The latter implies that all trajectories that start in $X_{I}$ avoid entering the unsafe set $X_{U}$, thus concluding the proof. ∎
To synthesise a certificate, we require complete knowledge of the behaviour $f$ of the dynamical system, to allow us to evaluate the derivative $\frac{dB}{dt}$. This may be impractical, and we therefore use data-driven techniques to learn a certificate.
III Data-Driven Certificate Synthesis
@@
We consider $N$ independent and identically distributed (i.i.d.) initial conditions, sampled according to $\mathbb{P}$, namely $\{x^{i}(0)\}_{i=1}^{N}\sim\mathbb{P}^{N}$. Initializing the dynamics from each of these initial states, we unravel a set of continuous-time trajectories $\{\xi^{i}\}_{i=1}^{N}$, also referred to as a multi-sample. Since there is no stochasticity in the dynamics, we can equivalently say that trajectories (generated from the random initial conditions) are distributed according to the same probabilistic law; hence, with a slight abuse of notation, we write $\xi\sim\mathbb{P}$. We impose the following assumption, standard in the scenario approach [16]. Intuitively, this assumption rules out degenerate problem instances, where we could select the same sample multiple times with nonzero probability.
Assumption 2 (Non-concentrated Mass)
@@
Since we are now dealing with a sample-based problem, we construct probabilistic certificates and hence probabilistic guarantees on the satisfaction of a given property.
III-A The Discrete-Time Case
Designing safety certificates that enjoy probabilistic guarantees when it comes to new unseen trajectories has been recently considered in [3], however, for discrete-time systems. We review the main developments in this direction, extending these to continuous-time ones. Consider a mapping $\mathcal{A}$ such that $B_{N}=\mathcal{A}(\{\xi^{i}\}_{i=1}^{N})$ as an algorithm that, based on $N$ samples, returns a certificate $B_{N}$. We call a compression set of such an algorithm any subset of the input multi-sample that returns the same certificate. That is, a sample subset $\mathcal{C}_{N}\subseteq\{\xi^{i}\}_{i=1}^{N}$ is a compression set if $\mathcal{A}(\mathcal{C}_{N})=\mathcal{A}(\{\xi^{i}\}_{i=1}^{N})$. In Algorithm 1, we provide a specific synthesis procedure through which $\mathcal{A}$ (and hence the certificate $B_{N}$) can be constructed. This algorithm is extended from [3] to continuous-time systems and is discussed in the next section.
Theorem 1 (Probabilistic Guarantees [3])
Consider Assumption 2, and let $B_{N}$ and $\mathcal{C}_{N}$ be the certificate and compression set, respectively, returned by Algorithm 1. Fix $\beta\in(0,1)$, and for $k<N$, let $\varepsilon(k,\beta,N)$ be the (unique) solution to the polynomial equation in the interval $[k/N,1]$
@@ III-A The Discrete-Time Case
Remark 1
Theorem 1 is not applicable to continuous-time trajectories. In particular, the inner probability refers to the probability of a new time-discretised trajectory $\tilde{\xi}$ being (un)safe. As a result, states along the state trajectory $\xi$ of the continuous-time system under study may violate the safety property (chiefly, between sampled states $x(t_{i})$ and $x(t_{i+1})$).
III-B Extension to the Continuous-Time Case
[…]
@@ III-B Extension to the Continuous-Time Case
$<\frac{1}{T}\Big{(}\inf_{x\in X_{U}}B(x)-\sup_{x\in X_{I}}B(x)\Big{)}-d,$
where $d\in\mathbb{R}$ is a tightening parameter. Define by $\mathcal{L}_{B}$ and $\mathcal{L}_{f}$ the Lipschitz constants of the certificate derivative $\frac{\partial B_{\theta^{\star}}}{\partial x}$ and of the dynamics $f(x)$ respectively, by $\mathcal{M}_{B},\mathcal{M}_{f}$ bounds on their norms, i.e. $\sup_{x}\|\frac{\partial B_{\theta^{\star}}}{\partial x}\|\leq\mathcal{M}_{B}$ and $\sup_{x}\|f(x)\|\leq\mathcal{M}_{f}$, and for $\overline{t}=\max_{k=1,\dots,M}(t_{k}-t_{k-1})$, define $d$ as
$d=\overline{t}\mathcal{M}_{f}\left(\mathcal{M}_{B}\mathcal{L}_{f}+\mathcal{M}_{f}\mathcal{L}_{B}\right).$ (12)
Theorem 2 (Continuous-Time Guarantees)
@@
The proof of this is achieved by bounding the difference between the safety of the continuous-time trajectories and their time-discretised approximations; see Appendix.
IV Certificate Synthesis
To learn a barrier certificate from samples, we consider a neural network, a well-studied class of function approximators that generalise well to a given task, although any other function approximation mechanism could be chosen instead. We consider a network with $N_{\mathrm{layers}}$ layers, and $N_{\mathrm{neurons}}$ neurons per layer. Denote all tunable neural network parameters by a vector $\theta$ $\in\mathbb{R}^{N_{\mathrm{layers}}N_{\mathrm{neurons}}}$. We then have that our certificate $B_{N}$ depends on $\theta$. For the results of this section, we write $B_{\theta}$ and drop the dependency on $N$ to ease notation.